110 likes | 191 Views
Summary of updated optics and PSL issues addressed during the LIGO Scientific Collaboration Meeting in August 2001, focusing on frequency noise reduction, improved optical components, and servo stabilization techniques.
E N D
Status of the LHO PSLs Rick Savage LIGO Hanford Observatory LIGO Scientific Collaboration MeetingAugust 13-16, 2001
Improved optics layout Optical table support legs NPRO frequency glitches Reference cavity tidal correction servo Intensity stabilization (Selected) PSL issues LSC Meeting - August 2001
Revised PSL/IOO optics layout LSC Meeting - August 2001
Improved frequency noise • Number of optical components reduced • Path lengths shortened • Reference cavity orientation changed LSC Meeting - August 2001
Frequency noise – table leg resonances Bolts tight Bolts loose Stacis system did not work-electronics noise We will try PEPS system ASAP LSC Meeting - August 2001
5 MHz 2 sec. NPRO frequency glitches • Approximately one per day • Alternates between step up and step down • T. Kane attributes to mode instability in pump laser diode LSC Meeting - August 2001
Tidal corection servo Reduce susceptibility to room temperature fluctuations 1 mK >> 1 mmcommon mode arm length LSC Meeting - August 2001
Tidal servo (cont.) LSC Meeting - August 2001
Tidal servo sensor noise 25 mK LSC Meeting - August 2001
Intensity stabilization servo • First loop of prototype system operational on2-km interferometer.(R. Karwoski, P. Russell, D. Ottaway) Inside the loopafter PMC Outside the loop after MC LSC Meeting - August 2001
RIN after modecleaner Only “inside” loopactive DP/P ~6.8e-8/rtHz Detector noise floor(laser light blocked) Light on PD = 0.5 mW >> 4.6e-8/rtHz shot noise limit LSC Meeting - August 2001